# Locking-Free Training of Physics-Informed Neural Network for Solving Nearly Incompressible Elasticity Equations

### v1

@@ 3 Robust PINNs for nearly incompressible equations
Using PINNs for solving the elasticity problem (1.1)-(1.2) for the case of nearly incompressible materials leads to a non-robust solution (locking), and the training of the PINN completely fails (see, e.g., Figure 1). As discussed earlier, this phenomenon is mainly due to the extreme imbalance between the coefficients $\mu$ and $\lambda$. To avoid the non-robustness issue of PINNs, we separate the effects of these coefficients by decomposing the given equations so that we can efficiently handle the large parameter $\lambda$. To be more specific, if we set ${\bf u}=\widehat{{\bf u}}+\widetilde{{\bf u}}$ for some $\widehat{{\bf u}}$ and $\widetilde{{\bf u}}$ in suitable function spaces, we may rewrite (1.1)-(1.2) as
$-\nabla\cdot\Big{(}\lambda\big{(}\nabla\cdot\big{(}\widehat{{\bf u}}+\widetilde{{\bf u}}\big{)}\big{)}{\bf I}\Big{)}-\nabla\cdot\Big{(}2\mu{\boldsymbol{\varepsilon}}(\widehat{{\bf u}}+\widetilde{{\bf u}})\Big{)}$ $={\bf f}\quad\,\text{in $\Omega$},$ (3.1)
$\widehat{{\bf u}}+\widetilde{{\bf u}}$ $={\bf{g}}\quad\text{on $\Gamma$}.$
@@
$-\nabla\cdot\Big{(}\lambda\big{(}\nabla\cdot\widehat{{\bf u}}\big{)}{\bf I}\Big{)}$ $=\widehat{{\bf f}}\quad\,\,\,\,\,\text{in $\Omega$},$ (3.2)
$-\nabla\cdot\Big{(}\mu{\boldsymbol{\varepsilon}}(\widehat{{\bf u}})\Big{)}$ $=\boldsymbol{0}\quad\,\,\,\,\text{in $\Omega$},$
@@
$-\nabla\cdot\Big{(}\lambda\big{(}\nabla\cdot\widetilde{{\bf u}}\big{)}{\bf I}\Big{)}$ $=\boldsymbol{0}\quad\quad\,\,\,\,\,\text{in $\Omega$},$ (3.3)
$-\nabla\cdot\Big{(}2\mu{\boldsymbol{\varepsilon}}(\widetilde{{\bf u}})\Big{)}$ $={\bf f}-\widehat{{\bf f}}\quad\,\text{in $\Omega$},$
$\widetilde{{\bf u}}$ $={\bf{g}}-\widehat{\bf{g}}\,\,\,\,\,\text{on $\Gamma$},$
for some intermediate forcing $\widehat{{\bf f}}$ and boundary condition $\widehat{{\bf{g}}}$. As we can see from the decomposed system (3.2)-(3.3), the original forcing term ${\bf f}$ is separated into two terms: $\widehat{{\bf f}}$ is the term related to $\lambda$, and ${\bf f}-\widehat{{\bf f}}$ is the term concerning $\mu$. By considering the terms associated with $\lambda$ and $\mu$ separately, we can expect that the imbalance between $\lambda$ and $\mu$ that causes locking will be resolved, and the negative impact of the magnitude of $\lambda$ for computing solutions is relatively alleviated.
Note here that the functions $\widehat{{\bf f}}$ and $\widehat{\bf{g}}$ are not known a priori. Therefore, we must solve the forward and inverse problems simultaneously in order to identify $\widehat{{\bf f}}$ and $\widehat{\bf{g}}$ that are well-suited to our setting. Since the above decomposition forms a coupled system of forward-inverse problems, the solutions of (3.2)-(3.3) may not be uniquely determined, as the system might become ill-posed. Noting that, under certain assumptions on the boundary $\Gamma$ of the body $\Omega$, and for a sufficiently regular $\widehat{\bf{g}}$, it is known from the (strongly elliptic) PDE theory that (3.2)2-(3.2)3 has a unique solution $\widehat{{\bf u}}(\widehat{\bf{g}})$ in the Sobolev space $[H^{2}(\Omega)]^{d}$. Then we can plug this solution into (3.2)1 to obtain $\widehat{{\bf f}}(\widehat{\bf{g}})$. We substitute this $\widehat{{\bf f}}(\widehat{\bf{g}})$ into (3.3)2 and solve (3.3)2-(3.3)3 to determine a unique solution $\widetilde{{\bf u}}(\widehat{\bf{g}})$ in $[H^{2}(\Omega)]^{d}$. If $\widetilde{{\bf u}}(\widehat{\bf{g}})$ also satisfies (3.3)1, we have found a solution that satisfies the entire coupled system. However, there is no guarantee that $\widetilde{{\bf u}}(\widehat{\bf{g}})$ is a solution of (3.3)1. Accordingly, we shall explore an alternative approach that minimizes the left-hand side of (3.3)1. More precisely, we plug the aforementioned $\widetilde{{\bf u}}(\widehat{\bf{g}})$ into (3.3)1, and consider the minimization problem:
${\bf{g}}_{\min}:=\operatorname*{arg\,min}_{\widehat{\bf{g}}}\Big{\|}\nabla\cdot\Big{(}\lambda\big{(}\nabla\cdot\widetilde{{\bf u}}(\widehat{\bf{g}})\big{)}{\bf I}\Big{)}\Big{\|},$ (3.4)
@@
Figure 3: A schematic diagram of the proposed method using a decomposition. Three independent neural networks are utilized for the three intermediate functions $\widehat{{\bf u}}$, $\widetilde{{\bf u}}$ and $\widehat{{\bf f}}$, and the loss functions are defined based on the decomposition (3.2)-(3.3).
@@
We shall now employ PINNs to solve the decomposed systems (3.2)–(3.3). Owing to the flexibility of PINNs in incorporating problem-specific physics through the design of appropriate loss functions, this framework enables a direct mathematical formulation of the previously discussed scenario, including the minimization problem (3.4). Moreover, although it is generally difficult to determine the minimizer ${\bf{g}}_{\min}$ of (3.4) and to set the boundary condition ${\bf{g}}$ and the intermediate forcing $\widehat{{\bf f}}$ in the decomposition (3.2)-(3.3), we will design the loss function for PINN to solve the forward-inverse problem, so that these functions are automatically adjusted appropriately during the training process of PINN. For the input value ${\boldsymbol{x}}$, we shall consider three independent fully-connected neural networks, where each neural network generates $\widehat{{\bf u}}$, $\widetilde{{\bf u}}$ and $\widehat{{\bf f}}$ as outputs respectively. At first, we initialize the parameters of the network for $\widehat{{\bf f}}$ so that $\widehat{{\bf f}}=\frac{1}{2}{\bf f}$ is satisfied. We then apply automatic differentiation to these output values in accordance with the systems in (3.2) and (3.3). Specifically, for random samples from the interior of the domain $\Omega$, we will compute the residuals of the equations in (3.2) and (3.3), and for random samples at the boundary, we will set up the loss function so that the sum of $\widehat{{\bf u}}$ and $\widetilde{{\bf u}}$ equals ${\bf{g}}$ so that the intermediate boundary value $\widehat{\bf{g}}$ is automatically adjusted during the training. The above discussions lead us to consider the following form of the loss function
$\mathcal{L}\Big{(}\widehat{{\bf u}},\widetilde{{\bf u}},\widehat{{\bf f}}\Big{)}=$ $\frac{\delta_{r}}{N_{r}}\sum^{N_{r}}_{i=1}\left(\Big{|}\nabla\cdot(\lambda\nabla\cdot\widehat{{\bf u}}\,{\bf I})({\boldsymbol{x}}^{i}_{r})+\widehat{{\bf f}}({\boldsymbol{x}}^{i}_{r})\Big{|}^{2}+\Big{|}\nabla\cdot(\mu{\boldsymbol{\varepsilon}}(\widehat{{\bf u}}))({\boldsymbol{x}}^{i}_{r})\Big{|}^{2}\right)$ (3.5)
@@
$+\frac{\delta_{s}}{N_{s}}\sum^{N_{s}}_{j=1}\left(\Big{|}\nabla\cdot(\lambda\nabla\cdot\widetilde{{\bf u}}\,{\bf I})({\boldsymbol{x}}^{j}_{s})\Big{|}^{2}+\Big{|}\nabla\cdot(2\mu{\boldsymbol{\varepsilon}}(\widetilde{{\bf u}}))({\boldsymbol{x}}^{j}_{s})+{\bf f}({\boldsymbol{x}}^{j}_{s})-\widehat{{\bf f}}({\boldsymbol{x}}^{j}_{s})\Big{|}^{2}\right)$
$+\frac{\delta_{t}}{N_{t}}\sum^{N_{t}}_{k=1}\Big{|}\widehat{{\bf u}}({\boldsymbol{x}}^{k}_{t})+\widetilde{{\bf u}}({\boldsymbol{x}}^{k}_{t})-{\bf{g}}({\boldsymbol{x}}^{k}_{t})\Big{|}^{2},$
@@
for some weight coefficients $\delta_{r}$, $\delta_{s}$ and $\delta_{t}$, where the interior points ${\boldsymbol{x}}^{i}_{r}$ and ${\boldsymbol{x}}^{j}_{s}$ are i.i.d. randomly sampled from the uniform distribution $\mathcal{U}(\Omega)$ and the boundary samples ${\boldsymbol{x}}^{k}_{t}$ are i.i.d. randomly chosen from $\mathcal{U}(\Gamma)$. The graphical illustration of the proposed method can be found in Figure 3. In the following section, we demonstrate through various numerical experiments that the proposed method is effective for the training of PINN for this problem and successfully resolves the locking phenomenon.

### v2

@@ 4 Robust PINNs for nearly incompressible equations
Using PINNs for solving the elasticity problem (1.1)-(1.2) for the case of nearly incompressible materials leads to a non-robust solution (locking), and the training of the PINN completely fails (see, e.g., Figure 1). As discussed earlier, this phenomenon is mainly due to the extreme imbalance between the coefficients $\mu$ and $\lambda$. To avoid the non-robustness issue of PINNs, we separate the effects of these coefficients by decomposing the given equations so that we can efficiently handle the large parameter $\lambda$. To be more specific, if we set ${\bf u}=\widehat{{\bf u}}+\widetilde{{\bf u}}$ for some $\widehat{{\bf u}}$ and $\widetilde{{\bf u}}$ in suitable function spaces, we may rewrite (1.1)-(1.2) as: with $\boldsymbol{\sigma}_{\lambda}({\bf v})=\lambda\,\bigl(\nabla\cdot{\bf v}\bigr)\,{\bf I}$ and $\boldsymbol{\sigma}_{\mu}({\bf v})=2\mu\,{\boldsymbol{\varepsilon}}({\bf v}),$
[…]
${\bf u}=\widehat{{\bf u}}+\widetilde{{\bf u}}$ $={\bf{g}}\quad\text{on $\Gamma$}.$
@@
[…]
@@
[…]
$\widetilde{{\bf u}}$ $=\widetilde{{\bf{g}}}\quad\,\,\,\,\,\,\text{on $\Gamma$},$
with
${\bf f}=\widehat{{\bf f}}_{\lambda}+\widetilde{{\bf f}}_{\mu}+\widehat{{\bf f}}_{\mu}\quad{\rm and}\quad{\bf{g}}=\widehat{\bf{g}}+\widetilde{\bf{g}}$. (4.4)
[…]
$\widehat{{\bf u}}_{2}({\boldsymbol{x}})=\widehat{{\bf u}}({\boldsymbol{x}})+Q{\boldsymbol{x}}+{\boldsymbol{d}}\quad{\rm and}\quad\widetilde{{\bf u}}_{2}({\boldsymbol{x}})=\widetilde{{\bf u}}({\boldsymbol{x}})-Q{\boldsymbol{x}}-{\boldsymbol{d}},$
where ${\boldsymbol{d}}\in\mathbb{R}^{d}$ and $Q$ is a skew-symmetric matrix, then $(\widehat{{\bf u}}_{2},\widetilde{{\bf u}}_{2})$ is also a solution of (4.2) and (4.3). More generally, choosing $\widetilde{{\bf u}}$ to satisfy (4.3)1 and (4.3)3 determines $\widetilde{{\bf f}}_{\mu}$ via (4.3)2, after which setting $\widehat{{\bf u}}:={\bf u}-\widetilde{{\bf u}}$ enforces (4.2). Thus, multiple load-boundary configurations can realize the same overall solution through the decomposition.
@@
Figure 4: A schematic diagram of the proposed method using a decomposition. Three independent neural networks are utilized for the three intermediate functions $\widehat{{\bf u}}$, $\widetilde{{\bf u}}$, and $\widehat{{\bf f}}$, and the loss functions are defined based on the decomposition (4.2)-(4.3).
@@
We shall now employ PINNs to solve the decomposed systems (4.2)-(4.3). Owing to the flexibility of PINNs in incorporating problem-specific physics through the design of appropriate loss functions, this framework enables a direct mathematical formulation of the previously discussed scenario. The auxiliary functions $\widehat{{\bf f}}_{\lambda},\widehat{{\bf f}}_{\mu},\widetilde{{\bf f}}_{\mu}$ as well as the boundary functions $\widehat{{\bf{g}}},\widetilde{{\bf{g}}}$ will themselves be modeled by neural networks, which will be automatically adjusted during the PINN training process. For the input value $\boldsymbol{x}$}, we shall consider four independent fully-connected neural networks, where each neural network generates $\widehat{{\bf u}}$, $\widetilde{{\bf u}}$, and $\widehat{{\bf f}}_{\lambda},\widehat{{\bf f}}_{\mu}$ as outputs, respectively. At first, we initialize the parameters of the network for $\widehat{{\bf f}}_{\lambda}$, $\widehat{{\bf f}}_{\mu}$ so that $\widehat{{\bf f}}_{\lambda}=\frac{1}{3}{\bf f}$ and $\widehat{{\bf f}}_{\mu}=\frac{1}{3}{\bf f}$ are satisfied. Note that our method’s performance is not sensitive to the choice $\frac{1}{3}$ here, and other values are also admissible. We then apply automatic differentiation to these output values in accordance with the systems in (4.2) and (4.3). Specifically, for random samples from the interior of the domain $\Omega$, we will compute the residuals of the equations in (4.2) and (4.3), and for random samples at the boundary, we will set up the loss function so that the sum of $\widehat{{\bf u}}$ and $\widetilde{{\bf u}}$ equals $\bf{g}$} so that the intermediate boundary value $\widehat{\bf{g}}$ is automatically adjusted during the training.
The above discussions lead us to consider the following form of the loss function
[…]
@@
$+\frac{1}{N_{s}}\sum^{N_{s}}_{j=1}\left(\delta_{s,1}\big|\nabla\cdot\boldsymbol{\sigma}_{\lambda}(\widetilde{{\bf u}})({\boldsymbol{x}}^{j}_{s})\big|^{2}+\delta_{s,2}\big|\nabla\cdot\boldsymbol{\sigma}_{\mu}(\widetilde{{\bf u}})({\boldsymbol{x}}^{j}_{s})+{\bf f}({\boldsymbol{x}}^{j}_{s})-\widehat{{\bf f}}_{\lambda}({\boldsymbol{x}}^{j}_{s})-\widehat{{\bf f}}_{\mu}({\boldsymbol{x}}^{j}_{s})\big|^{2}\right)$
$+\frac{\delta_{b}}{N_{b}}\sum^{N_{b}}_{k=1}\big|\widehat{{\bf u}}({\boldsymbol{x}}^{k}_{b})+\widetilde{{\bf u}}({\boldsymbol{x}}^{k}_{b})-{\bf{g}}({\boldsymbol{x}}^{k}_{b})\big|^{2},$
where for some weight parameters $\delta_{r},\delta_{s},\delta_{b}$, we set $\delta_{r,1}=\delta_{r,2}=\delta_{r}$ and $\lambda^{2}\delta_{s,1}=\delta_{s,2}=\delta_{s}$. In the case of inhomogeneous $\lambda$, where $\lambda({\boldsymbol{x}})=\Lambda\tilde{\lambda}({\boldsymbol{x}})$ for some constant $\Lambda$, we use $\delta_{s,1}=\delta_{s}/\Lambda^{2}$.
@@
The interior points ${\boldsymbol{x}}^{i}_{r}$ and ${\boldsymbol{x}}^{j}_{s}$ are i.i.d. randomly sampled from the uniform distribution $\mathcal{U}(\Omega)$ and the boundary samples ${\boldsymbol{x}}^{k}_{b}$ are i.i.d. randomly chosen from $\mathcal{U}(\Gamma)$. The graphical illustration of the proposed method can be found in Figure 4. In the following section, we demonstrate through various numerical experiments that the proposed method is effective for the training of PINN for this problem and successfully resolves the locking phenomenon.
5 Error analysis
In the following, we analyze the approximation error under the assumption that the optimization procedure yields a neural network which approximates the true solution. The inner product on the space ${\bf L}^{2}(\Omega)=[L^{2}(\Omega)]^{d}$ is denoted by $\langle\cdot,\cdot\rangle$, and is the associated norm is written as $\|\cdot\|$, where for ${\bf v},{\bf w}\in{\bf L}^{2}(\Omega)$,
$\langle{\bf v},{\bf w}\rangle=\int_{\Omega}{\bf v}\cdot{\bf w}\,{\rm{d}}{\boldsymbol{x}}~~{\rm and}~~\|{\bf v}\|=\sqrt{\langle{\bf v},{\bf v}\rangle}.$
For the inner product and the norm on ${\bf L}^{2}(\partial\Omega)$, we use the notations $\langle{\bf v},{\bf w}\rangle_{\partial\Omega}$ and $\|\cdot\|_{\partial\Omega}$.
Let ${\bf H}^{2}(\Omega)=[H^{2}(\Omega)]^{d}$, where $H^{2}(\Omega)$ is the standard Sobolev space. The norm on ${\bf H}^{2}(\Omega)$ is denoted by $\|\cdot\|_{2}.$ The error is measured using the fractional norm $\|\cdot\|_{1/2}$, defined by: $\|{\bf w}\|^{2}_{1/2}=\|{\bf w}\|^{2}+|{\bf w}|_{1/2}^{2}$, where $|\cdot|_{1/2}$ is the (Slobodeckij) seminorm on the fractional Sobolev space $H^{1/2}(\Omega)]^{d}$ defined as:
$|{\bf w}|^{2}_{1/2}=\int_{\Omega}\int_{\Omega}\frac{|{\bf w}({\boldsymbol{x}})-{\bf w}({\boldsymbol{y}})|^{2}}{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{d+1}}\,{\rm{d}}{\boldsymbol{x}}\,{\rm{d}}{\boldsymbol{y}}.$
Define the following energy functional ${\mathcal{E}}:{\bf H}^{2}(\Omega)\to\mathbb{R}$ by
${\mathcal{E}}({\bf v})=\delta_{r}\|\nabla\cdot\boldsymbol{\sigma}({\bf v})+{\bf f}\|^{2}+\delta_{b}\|{\bf v}-{\bf{g}}\|^{2}_{\partial\Omega},$ (5.1)
where $\delta_{r},\delta_{b}$ are constants as defined in (2.2). Solving the linear elasticity equation (1.1) is equivalent to minimizing the energy functional $\mathcal{E}$} over the Sobolev space ${\bf H}^{2}(\Omega)$.
A neural network with $L$ layers, each layer has $n_{\ell}$ neurons, for $\ell=0,\ldots,L$ can be modeled as the following structure
${\boldsymbol{\theta}}=(W_{1},{\bf b}_{1},W_{2},{\bf b}_{2},\ldots,W_{L},{\bf b}_{L}),$
where $W_{\ell}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}}$, for $\ell=1,2,\ldots,L$ are the weight matrices and ${\bf b}_{\ell}\in\mathbb{R}^{n_{\ell}}$ are the bias vectors. Given an input vector ${\bf z}_{0}={\boldsymbol{x}}\in\mathbb{R}^{n_{0}}$, the output vector ${\bf a}_{L}$ of the network $\boldsymbol{\theta}$} is defined by
${\bf a}_{L}=W_{L}{\bf z}_{L-1}+{\bf b}_{L},$ (5.2)
where ${\bf z}_{\ell}=\rho(W_{\ell}{\bf z}_{\ell-1}+{\bf b}_{\ell})$ for $\ell=1,2,\ldots,L-1,$ and $\rho$ is some given activation function. The neural network function with parameters $\boldsymbol{\theta}$} and the activation function $\rho$ is the function
${\bf v}_{{\boldsymbol{\theta}}}:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{L}},\quad{\boldsymbol{x}}\mapsto{\bf a}_{L}({\boldsymbol{x}}).$ (5.3)
We denote by $\Theta$ the parameter space which contains all possible values of $\boldsymbol{\theta}$}. It can be seen that $\Theta=\mathbb{R}^{p}$ for some integer $p$ which depends on the values $n_{\ell}$, for $\ell=0,1,\ldots,L$. The collection of all network functions with a certain parameter space $\Theta$ is denoted by ${\mathcal{F}}_{\Theta}$. The PINN approach to solve the linear elasticity equation (1.1) is to minimize the energy $\mathcal{E}$} given in (5.1) over the function class ${\mathcal{F}}_{\Theta}$. More precisely, the network will minimize a loss functional $\mathcal{L}:\Theta\to\mathbb{R}$ (see (2.2)), such that $\mathcal{L}({\boldsymbol{\theta}})\approx{\mathcal{E}}({\bf u}_{{\boldsymbol{\theta}}}).$ Let $\boldsymbol{\theta}$$^{*}$ minimize $\mathcal{L}({\boldsymbol{\theta}})$ inexactly and let ${\bf u}_{{\boldsymbol{\theta}}^{*}}$ be the corresponding neural network solution, then
$\|{\bf u}_{{\boldsymbol{\theta}}^{*}}-{\bf u}\|^{2}_{1/2}\lesssim\mathcal{L}({\boldsymbol{\theta}}^{*})+\eta({\boldsymbol{\theta}}^{*}),$ (5.4)
where $\eta({\boldsymbol{\theta}}^{*})={\mathcal{E}}({\bf u}_{{\boldsymbol{\theta}}^{*}})-\mathcal{L}({\boldsymbol{\theta}}^{*})$ is a quadrature error. Furthermore,
$\|{\bf u}_{{\boldsymbol{\theta}}^{*}}-{\bf u}\|^{2}_{1/2}\lesssim\kappa({\bf u}_{{\boldsymbol{\theta}}^{*}})+\inf_{{\bf u}_{{\boldsymbol{\theta}}}\in{\mathcal{F}}_{\Theta}}\|{\bf u}_{{\boldsymbol{\theta}}}-{\bf u}\|^{2}_{2},$ (5.5)
where the optimization error $\kappa({\bf u}_{{\boldsymbol{\theta}}^{*}})={\mathcal{E}}({\bf u}_{{\boldsymbol{\theta}}^{*}})-\inf_{{\bf u}_{{\boldsymbol{\theta}}}}{\mathcal{E}}({\bf u}_{{\boldsymbol{\theta}}})$. See [42] for the proof of these error estimate results. We aim next to show similar results for the proposed decomposed PINN method. For convenience, we introduce the following notations: let
$\boldsymbol{\sigma}_{D}=\begin{bmatrix}\boldsymbol{\sigma}_{\lambda}\\ \boldsymbol{\sigma}_{\mu}\end{bmatrix},\quad\widehat{\bf f}_{D}=\begin{bmatrix}\,\widehat{{\bf f}}_{\lambda}\,\\ \widehat{{\bf f}}_{\mu}\end{bmatrix},\quad\widetilde{\bf f}_{D}=\begin{bmatrix}\,\,{\bf 0}\,\\ \widetilde{{\bf f}}_{\mu}\,\end{bmatrix},\quad{\bf v}_{D}=\begin{bmatrix}\,\widehat{\bf v}\\ \widetilde{\bf v}\,\end{bmatrix},\quad{\bf v}_{S}=\widehat{\bf v}+\widetilde{\bf v}$.
The loss function $\mathcal{L}_{D}$ is defined as in (4.5) with $\delta_{r,1}=\delta_{r,2}=\delta_{r}$ and $\delta_{s,1}=\delta_{s,2}=\delta_{s}$. The associated energy functional
${\mathcal{E}}_{D}({\bf v}_{D})$ $=\delta_{r}\|\nabla\cdot\boldsymbol{\sigma}_{D}(\widehat{\bf v})+\widehat{{\bf f}}_{D}\|^{2}+\delta_{s}\|\nabla\cdot\boldsymbol{\sigma}_{D}(\widetilde{\bf v})+\widetilde{{\bf f}}_{D}\|^{2}+\delta_{b}\|{\bf v}_{S}-{\bf{g}}\|^{2}_{\partial\Omega},$ (5.6)
where it is understood that in the first two terms on the RHS of equation (5.6) the vectors are in $L^{2}(\Omega)]^{2d}$. Hence, our PINN approach is to solve two systems (4.2) and (4.3) simultaneously by minimizing ${\mathcal{E}}_{D}$ over the function class ${\mathcal{F}}_{\Theta}\times{\mathcal{F}}_{\Theta}$. More precisely, the networks will minimize a loss functional
$\mathcal{L}_{D}:\Theta\times\Theta\to\mathbb{R},\quad\mathcal{L}_{D}({\boldsymbol{\theta}}_{D})=\mathcal{L}({\boldsymbol{\theta}}_{D})\approx{\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}})={\mathcal{E}}_{D}(\widehat{\bf u}_{\theta_{1}},\widetilde{\bf u}_{\theta_{2}}).$
Define the following bilinear operator $a_{D}:[H^{2}(\Omega)]^{2d}\times[H^{2}(\Omega)]^{2d}\to\mathbb{R}$ as
$a_{D}({\bf v}_{D},{\bf w}_{D})=\delta_{r}\langle\nabla\cdot\boldsymbol{\sigma}_{D}(\widehat{\bf v}),\nabla\cdot\boldsymbol{\sigma}_{D}(\widehat{\bf w})\rangle+\delta_{s}\langle\nabla\cdot\boldsymbol{\sigma}_{D}(\widetilde{\bf v}),\nabla\cdot\boldsymbol{\sigma}_{D}(\widetilde{\bf w})\rangle+\delta_{b}\langle{\bf v}_{S},{\bf w}_{S}\rangle_{\partial\Omega}.$ (5.7)
Define the bilinear operator $a:{\bf H}^{2}(\Omega)\times{\bf H}^{2}(\Omega)\to\mathbb{R}$ as
$a({\bf v},{\bf w})=\delta_{r}\langle\nabla\cdot\boldsymbol{\sigma}({\bf v}),\nabla\cdot\boldsymbol{\sigma}({\bf w})\rangle+\delta_{b}\langle{\bf v},{\bf w}\rangle_{\partial\Omega},$ (5.8)
Hence, by linearity and with the help of the Cauchy–Schwarz inequality,
$a({\bf v}_{S},{\bf v}_{S})$ $=\delta_{r}\|\nabla\cdot(\boldsymbol{\sigma}_{\lambda}+\boldsymbol{\sigma}_{\mu})(\widehat{\bf v}+\widetilde{\bf v})\|^{2}+\delta_{b}\|{\bf v}_{S}\|^{2}_{\partial\Omega}$
$\lesssim\|\nabla\cdot\boldsymbol{\sigma}_{\lambda}(\widehat{\bf v})\|^{2}+\|\nabla\cdot\boldsymbol{\sigma}_{\mu}(\widehat{\bf v})\|^{2}+\|\nabla\cdot\boldsymbol{\sigma}_{\lambda}(\widetilde{\bf v})\|^{2}+\|\nabla\cdot\boldsymbol{\sigma}_{\mu}(\widetilde{\bf v})\|^{2}+\delta_{b}\|{\bf v}_{S}\|^{2}_{\partial\Omega}$
$\lesssim a_{D}({\bf v}_{D},{\bf v}_{D}).$
Therefore, by using the coercivity property of $a(\cdot,\cdot)$ in [42, Lemma 3], we deduce that for $\widehat{\bf v},\,\widetilde{\bf v}\in{\bf H}^{2}(\Omega)$, there holds
$\|{\bf v}_{S}\|^{2}_{1/2}=\|\widehat{\bf v}+\widetilde{\bf v}\|^{2}_{1/2}\lesssim a({\bf v}_{S},{\bf v}_{S})\lesssim a_{D}({\bf v}_{D},{\bf v}_{D}).$ (5.9)
Theorem 5.1.
Let ${\boldsymbol{\theta}}^{*}_{D}$ minimize $\mathcal{L}_{D}({\boldsymbol{\theta}}_{D})$ inexactly, and let ${\bf u}_{D,{\boldsymbol{\theta}}^{*}}$ be the corresponding neural network solutions, then
$\|{\bf u}_{S,{\boldsymbol{\theta}}^{*}}-{\bf u}\|^{2}_{1/2}\lesssim\mathcal{L}_{D}({\boldsymbol{\theta}}^{*}_{D})+\eta_{D}({\boldsymbol{\theta}}^{*}_{D}),$ (5.10)
where ${\bf u}_{S,{\boldsymbol{\theta}}^{*}}=\widehat{{\bf u}}_{\theta_{1}^{*}}+\widetilde{{\bf u}}_{\theta_{2}^{*}}$, ${\bf u}={\bf u}_{S}=\widehat{{\bf u}}+\widetilde{{\bf u}}$ and $\eta_{D}({\boldsymbol{\theta}}^{*}_{D})={\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}})-\mathcal{L}_{D}({\boldsymbol{\theta}}^{*}_{D})$ is a quadrature error. Furthermore,
$\|{\bf u}_{S,{\boldsymbol{\theta}}^{*}}-{\bf u}\|^{2}_{1/2}\lesssim\kappa({\bf u}_{D,{\boldsymbol{\theta}}^{*}})+\inf_{{\bf u}_{D,{\boldsymbol{\theta}}}\in{\mathcal{F}}_{\Theta}\times{\mathcal{F}}_{\Theta}}\|{\bf u}_{D,{\boldsymbol{\theta}}}-{\bf u}_{D}\|^{2}_{2},$ (5.11)
with $\kappa({\bf u}_{D,{\boldsymbol{\theta}}^{*}}):={\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}})-\inf_{{\bf u}_{D,{\boldsymbol{\theta}}}\in{\mathcal{F}}_{\Theta}\times{\mathcal{F}}_{\Theta}}{\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}})$.
Proof.
Let $\widehat{{\bf u}}$ and $\widetilde{{\bf u}}$ solve the systems (4.2) and (4.3). Then ${\mathcal{E}}_{D}({\bf u}_{D})=0$ and it follows from the definition of the bilinear form (5.7) that
${\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}})-{\mathcal{E}}_{D}({\bf u}_{D})=a_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}}-{\bf u}_{D},{\bf u}_{D,{\boldsymbol{\theta}}^{*}}-{\bf u}_{D}).$ (5.12)
Using the coercivity of the bilinear form $a_{D}(\cdot,\cdot)$, see (5.9), we obtain
$\|{\bf u}_{S,{\boldsymbol{\theta}}^{*}}-{\bf u}\|^{2}_{1/2}\lesssim{\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}})=\mathcal{L}_{D}({\boldsymbol{\theta}}_{D}^{*})+\eta_{D}({\boldsymbol{\theta}}^{*}_{D}),$ (5.13)
proving the estimate (5.10). To prove the second estimate (5.11), by using the definition of the bilinear form, we have, for any ${\bf u}_{D,{\boldsymbol{\theta}}}\in[H^{2}(\Omega)]^{2d}$,
${\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}})-{\mathcal{E}}_{D}({\bf u}_{D})=a_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}}-{\bf u}_{D},{\bf u}_{D,{\boldsymbol{\theta}}^{*}}-{\bf u}_{D})\lesssim\|{\bf u}_{D,{\boldsymbol{\theta}}}-{\bf u}_{D}\|^{2}_{2}.$
Now using the above estimate
${\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}})-{\mathcal{E}}_{D}({\bf u}_{D})$ $={\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}^{*}})-\inf_{\theta_{1},\theta_{2}}{\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}})+\inf_{\theta_{1},\theta_{2}}[{\mathcal{E}}_{D}({\bf u}_{D,{\boldsymbol{\theta}}})-{\mathcal{E}}_{D}({\bf u}_{D})]$ (5.14)
$\leq\kappa({\bf u}_{D,{\boldsymbol{\theta}}^{*}})+C\|{\bf u}_{D,{\boldsymbol{\theta}}}-{\bf u}_{D}\|^{2}_{2}.$
The second estimate now follows from (5.12), (5.13) and (5.14). ∎